(1 point) Consider the system of differential equations dx dt = -1.6x + 0.5y, dy dt = 2.5x – 3.6y. For this system, the small(1 point) Calculate the eigenvalues of this matrix: [Note-- youll probably want to use a calculator or computer to estimate(1 point) a. Find the most general real-valued solution to the linear system of differential equations x -8 -10 8 x. xi(t) =
Answer #1

du t 는 16ut osg =프리 아 글 5 dy y: g.su -30g t! U 18 anto g X= AX where x- ㅠ [5] A= g S (A-11 ܟ ܐ ,; A+에 CO 100 베영 요 5 =) LrL (As eigen values are Real ; unequal and of Same Sign (-ve) son Mode and Stable Hence Y=AY is is Diff. Diff egno then 5+ has S> Hit 1250 Ugal Vi= (4)=* for la = TO (A-()I)v=0 15) 4(67 10) Hi-Ug Hg = Vya EC!! eigen vector V=( کر on! (een Solution, Xitl> Put t-o yw ) Xole (4) = ) 15 (-)) + -> 9565 +1월 64 2 +) 키 او را در 스 2 (= 5 3 ) g Yt5 1구 3. 3 ㅋh ㅋ X() 의 lo e 17 이 3 + ( )